OpenStudy (minthia):

What transformations change the graph of f(x) to the graph g(x) f(x)=-7 x 2 g(x)=-35x2+5

OpenStudy (anonymous):

do u still need help :)

OpenStudy (minthia):

yes

OpenStudy (anonymous):

the parent function is x^2, it seems that x^2 is becoming (x+7)^2, this transformation is like F(x)=x^2 => F(X+7)=(x+7)^2 the result of this is shifting the graphs origin 7 units left The addition of 9 is an addition outside the parent function which shifts it up or down. therefore g(x) is a linear transform of 7 units left and 9 units up

OpenStudy (minthia):

wow thanks for explains it do you think you can help me with another one like this?

OpenStudy (anonymous):

sure, hit me with ur best shot. lol

OpenStudy (minthia):

what transformation change the graph f (X) =x2 g(x)= (x+3)2 - 7

OpenStudy (anonymous):

it is going to shift lift.

OpenStudy (anonymous):

no sorry wait

OpenStudy (anonymous):

shift up which is a Vertical shift

OpenStudy (jhannybean):

think of \((x+3)^2-7\) as \((x-(-3))^2-7\). When you look at it this way, you can clearly tell that the x-intercept is negative and to the let, therefore the graph shifts left by 3 units.

OpenStudy (jhannybean):

left*

OpenStudy (anonymous):

so I was right the first time @jhanny that it would shift left